OpenStudy (anonymous):

Convert 1539 to a numeral in base 7

OpenStudy (anonymous):

You have to use powers of 7: 7^0 = 1, 7^1 = 7, 7^2 = 49, 7^3 = 343. Start by seeing how many 343's go into 1539. You will get remainder and the number of 343's. That latter number is your "thousands". Take the remainder and see how many 49's you have in that number. That's your "hundreds". Take remainder and do 7's and then do the ones.

OpenStudy (anonymous):

Are you following this?